OpenStudy (anonymous):

Derived \[\sin (\frac{ x+1 }{ 2x-3 } )\]

OpenStudy (unklerhaukus):

do you remember the chain rule ?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

\[(f\circ g)'(t)=f'(g(t))\cdot g'(t)\]

OpenStudy (unklerhaukus):

where \(f\circ g=f(g(t))\)

OpenStudy (unklerhaukus):

in your question f =sin

OpenStudy (unklerhaukus):

can you tell me what g(x) is ?

OpenStudy (anonymous):

sin

OpenStudy (unklerhaukus):

nope,

OpenStudy (anonymous):

then .. can u solve this derived

OpenStudy (unklerhaukus):

To take the derivative of the function-of-a-function, you have to be able to identify the functions

OpenStudy (unklerhaukus):

your function is \[y=\sin\left(\frac{x+1}{2x−3}\right)\] you need to decompose \(y=f(g(x))\) \[f=\sin \\\, \\\,\\ g(x)=?\]

OpenStudy (anonymous):

x+1/2x-3

OpenStudy (unklerhaukus):

ok, so now look at the rule for takig the derivative of a function-of-a-function \[(f(g(t)))'=(f∘g)'(t)=f'(g(t))⋅g'(t)\]

OpenStudy (unklerhaukus):

you need to take the derivative of f, \[f=\sin\\f'= \]

OpenStudy (anonymous):

cos

OpenStudy (unklerhaukus):

correct ,

OpenStudy (unklerhaukus):

now all you need before you can subsitut into that formula is to find \[g'(t)\] this is a bit tricker , and you might want to use the quotient rule \[y=\frac{u}{v}\\y'=\frac{vu'-uv'}{v^2}\]

OpenStudy (unklerhaukus):

so you will have to decompose \[\frac{x+1}{2x−3}\] onto \(\dfrac uv\) form and then find \(u,u',v,v',v^2\)

OpenStudy (anonymous):

nd cos ..?

OpenStudy (unklerhaukus):

pardon

OpenStudy (anonymous):

no pasa nada

OpenStudy (anonymous):

np prob

OpenStudy (unklerhaukus):

do you still need any help on this question?

OpenStudy (anonymous):

yes